# Representing Molecules with Algebraic Data Types: Beyond SMILES and SELFIES

### v4

@@ Abstract
Algebraic data types (ADTs) let a representation specify—at the type level—what molecular values are valid and what transformations are meaningful. We propose a molecular representation as a family of typed ADTs that separates (i) constitution (Dietz-style bonding systems), (ii) 3D coordinates and stereochemistry, and (iii) electronic-structure annotations. This separation makes invariants explicit, supports deterministic local edits, and provides hooks for symmetry-aware and Bayesian modeling. These data structures, allow us to consider how the representation constrains operations which may be performed over them. Types make invalid manipulations unrepresentable and make it easier to define meaningful priors/likelihoods over generative models (programs with sample and score operations). Unlike string-based formats, the ADT exposes chemical structure directly; validity conditions (e.g., valence and symmetry constraints) can be enforced by construction and checked deterministically during transformations. We optionally attach electronic-structure annotations (shell/subshell/orbital metadata) to atoms when such information is available; we do not attempt to compute orbitals in this work.
We sketch Bayesian probabilistic programming via an integration with LazyPPL, a lazy probabilistic programming library; molecules can be made instances of a group under rotation to support geometric learning settings where molecular properties are invariant under rigid motions and relabellings;; and the framework’s flexibility is demonstrated through an extension to represent chemical reactions.
We provide a Haskell library implementing the representation, released under an OSI-approved open-source license and archived with a DOI.
Scientific contribution: We introduce a typed molecular data model based on ADTs with a Haskell reference implementation. The model combines a Dietz-style constitution layer with geometry and electronic annotations, and it is designed to support validation, deterministic transformations, and integration into Bayesian and geometric learning workflows.
@@ 1 Introduction
The choice of molecular representation affects cheminformatics workflows such as molecular property prediction and generative design [1]. More expressive representations can make it easier to encode the information required by downstream models, but performance also depends on computational cost, ease of integration, and how reliably the representation supports valid edits, comparison, and conversion [2, 3]. Here we explore an algebraic approach: representing molecules as typed values built from constructors for atoms, connectivity, stereochemistry, and geometry, together with composable operations on that structure.
Before continuing with the article, we introduce the terms of art used throughout. These informal definitions are tuned to our setting; for formal statements see [4, 5].
Definition 1 (Data types and algebraic data types (ADTs)).
A data type is a compile-time contract specifying (i) which values exist and (ii) which operations are meaningful on those values. An algebraic data type (ADT) defines values by a finite family of constructors (products/records and sums/variants, possibly recursive) and is deconstructed by pattern matching. In this work, we represent molecules as ADTs so that parsing, validation, and transformation operate on structured molecular objects (atoms, bonds, bonding systems, geometry, and annotations) rather than on ad hoc string edits.
@@
Despite decades of work, string encodings such as SMILES [6] and, more recently, SELFIES [7] remain widely used interchange and model-input formats [8, 9]. Their convenience for storage and exchange comes with well-known limitations: standard SMILES is primarily graph-level and cannot directly represent many complex bonding patterns (e.g. general delocalisation or multicentre bonding); strings do not include explicit 3D information (coordinates, torsions), and stereochemical coverage is uneven across edge cases. In addition, SMILES is syntactically brittle – models can emit invalid strings—and small token edits can induce large structural changes, complicating learning and property prediction [7]. Finally, token sequences do not expose symmetry actions (e.g. $E(3)$) or typed structure directly, so geometric or Bayesian workflows typically require additional parsing and featurisation steps.
Fingerprints encode molecules into fixed-length feature vectors for retrieval, clustering, and QSAR. They are effective as descriptors, but can be lossy with respect to stereochemistry and geometry and are not, by themselves, canonical molecular representations. In this paper we therefore treat fingerprints as downstream descriptors derived from a structured representation, not as the representation itself. These approaches may be effective for retrieval and similarity ranking, but they can be lossy with respect to stereochemistry or geometry, and they often require careful choices of similarity metrics and calibration when used for learning or uncertainty estimation [10, 11, 12]. In this paper we treat fingerprints as downstream descriptors rather than as a primary molecular representation.
Krenn et al. also proposed a “molecular programming language”, noting that strings can encode Turing‑complete languages [1]. That observation is correct, but not prescriptive. Rather than extending a string grammar until it behaves like a language, one can treat the molecule itself as a first‑class program object in an ordinary programming language, with explicit types and operations defined on it. In a functional paradigm, ADTs paired with pure functions may provide composability, equational reasoning about transformations on molecules, and immutability, helping to make invariants and exceptions explicit. We propose that if molecules are to be generated, validated, optimised and transformed by programs, representing them directly as typed data – rather than as tokens that must be decoded – offers a more transparent and verifiable route.
@@
A brief terminological note: conflating a storage or transmission format, e.g. whether it is transmitted as a string on Wikipedia (e.g., a SMILES or InChI string) or as an SDF file with a computational data type (the in‑memory structure and its operations) has led to methodological confusion. We argue that this conflation has nudged cheminformaticians toward grammar engineering and away from semantically faithful, geometry‑aware modeling. The algebraic view places data types first: molecules as typed values with constructors for atoms, bonds, stereochemistry, and 3D information (including dihedral angles), supporting modelling choices where invariances, validity, and explicit structure matter (e.g., geometric deep learning and uncertainty-aware inference).
Other work has linked functional programming and cheminformatics before [13, 14], though this work was primarily focused on implementing existing cheminformatics workflows within functional programming languages, and the benefits of immutable state and type‑safety. Instead, we provide a molecular representation and the surrounding context for the importance of considering data types when choosing a representation.
@@
1.1 Molecular Graphs and the Dietz representation
Multigraphs retain familiar binary bonds while explicitly encoding delocalised, multi-centre electron sharing without proliferating ad hoc bond types [15].
@@ 1.1.1 Molecular Graphs
Many molecular representations are built on molecular graphs and thus inherit their strengths and limitations [16]. String encodings such as SMILES and SELFIES decode to molecular graphs, and SDF molfiles organise “blocks” for atoms and bonds.
Definition 2 (Graph).
A graph $G=(V,E)$ is a pair consisting of a set of vertices $V$ and an edge set $E\subseteq V\times V$. [^1] Set-theoretic notation used once: $V$ is the vertex set and $E\subseteq V\times V$ the set of ordered pairs $(u,v)$; for undirected graphs $(u,v)$ and $(v,u)$ are treated equivalently.
A molecular graph has vertices as atoms and edges as bonds; such graphs are connected, undirected, and labeled. Vertex labels may include element and charge; edge labels may include bond type/order or stereochemical annotations [15, 16].
@@ 1.1.3 Multigraphs
A multigraph $G=(V,E)$ has vertex set $V$ and a multiset (bag) $E$ of edges—edges may repeat. Multigraphs appear naturally when multiple bonding systems contribute between the same atom pair.
An edge in the multigraph corresponds to a bonding relation between a specific atom pair $\{i,j\}\in F$ and is labeled by the identifier of its bonding system and the shared electron count $s$. Localised bonds are the special case $|F|=1$; delocalised or multi-centre bonding is captured by a single bonding system whose $F$ contains multiple atom pairs (e.g., the benzene $\pi$ sextet spans six C–C pairs). Multiple bonding systems may contribute between the same two atoms (e.g., $\sigma$ and $\pi$ components), which is why a multigraph—rather than a simple graph—is required [15].
@@ 1.1.4 Dietz representation
The Dietz representation provides a computational framework capable of representing any chemical structure that can be described by a structural formula, overcoming the limitations of (non-multi)graph-based and hypergraph-based methods (including string encodings that decode to molecular graphs) [15]. Dietz factors structure into constitution, configuration, and conformation. Here we focus on constitution, since our ADT supplies 3D coordinates separately.
Definition 3 (Constitution).
@@
Induced labeled multigraph (from $C=(V,B)$).
Let $E^{\ast}=\{\,(\{i,j\},b)\mid b=(s,F)\in B,\ \{i,j\}\in F\,\}$. The induced multigraph is $G^{\ast}=(V,E^{\ast})$ with edge-labeling $\ell\colon(\{i,j\},b)\mapsto(\operatorname{id}(b),s)$. This makes explicit how $(V,B)$ induces a labeled molecular multigraph while preserving pairwise adjacencies and capturing delocalisation via edge-induced subgraphs labeled by bonding-system identifiers.
Benzene can be visualised as Lewis Structure as follows:
@@
Equivalently, the constitution can be viewed as a multigraph together with metadata grouping edges into electron-sharing systems.
A major advantage of the Dietz representation is its general model for the representation of bonding. As a result, it is straightforward to describe complex bonding systems such as organometallics, with ferrocene and diborane being given as examples in the original article. The lack of labeling requirements for delocalised bonding stands in contrast to string-based representations such as SMILES, which has explicit rules in the grammar (a form of labeling) for aromatic carbon, and two rules for different cases of aromatic nitrogen.
@@
Figure 2: Three‑dimensional view of ferrocene showing the metallocene “sandwich” geometry. Image Credit: [19].
@@
Despite satisfying the three requirements identified in [1] for representation beyond organic chemistry—namely, the ability to represent delocalized bonding, explicit labeling of terminal hydrogen atoms, and advanced stereochemistry such as in cumulenes—Krenn et al. describe the Dietz representation (apparently mischaracterized as using hypergraphs) as one that “leads to complicated nested sets of brackets that may be hard to comprehend” [1].
Although set-theoretic notation may be unfamiliar, the definition or “grammar” of the constitution of a molecule can be efficiently described. This stands in contrast to SMILES strings or SELFIES strings (which as an additional challenge contains overloading of symbols in the grammar [1]).
@@ 1.2.1 SMILES
SMILES encodes molecular graphs as ASCII strings using atoms, bonds, branches, and ring closures, with optional isomeric annotations for stereochemistry. As a line notation, SMILES is compact, widely used [16, 17, 1, 20, 21, 22, 23] and widely interoperable; however, it does not store three‑dimensional coordinates or conformers and, like any graph‑level encoding, it must rely on external geometry to recover dihedrals and conformations.[24] In practice, limitations arise around (i) resonance/tautomerism and delocalisation (which require multiple structures or conventions), (ii) the breadth of stereochemical cases that can be expressed at the graph level, and (iii) variation across toolchains (e.g., aromaticity perception, canonicalisation).[17] These are active areas of improvement: OpenSMILES provides a community spec, and newer proposals such as BALSA offer a formally specified subset with explicit syntax/semantics to reduce ambiguity across implementations [25].
@@
SMILES representations are non-unique: a single molecule can be expressed by multiple valid strings [6]. As well as complicating database searches [17], machine learning algorithms must discern the equality of different representations. Although some machine learning approaches have attempted to utilize multiple SMILES for the same molecule in their training sets, and have claimed improvements to their models as a result [32, 33], these gains may reflect regularisation effects from representing the same graph with multiple strings, but they also underline that the learning objective must spend capacity on the string grammar rather than chemistry. This motivates representations whose locality and invariants are explicit.. In other words, it exposes a weakness of SMILES for machine learning, not an advantage.
SMILES does not explicitly represent molecular symmetries (graph automorphisms) as first-class objects. For example, the Cyclopentadienyl anion in Figure 3 can be written as [cH-]1cccc1, but the rotational symmetry evident in Fig. 1 is not encoded in the string. Moreover, because SMILES is non-unique, alternative strings such as 1cccc1[cH-] represent the same structure without any explicit indication of equivalence at the representation level.

### v5

@@ Abstract
Benchmarks of molecular machine learning models often treat the molecular representation as a neutral input format, yet the representation defines the syntax of validity, edit operations, and invariances that models implicitly learn. We propose MolADT, a typed intermediate representation (IR) for molecules expressed as a family of algebraic data types that separates (i) constitution via Dietz-style bonding systems, (ii) 3D geometry and stereochemistry, and (iii) optional electronic annotations. By shifting from string edits to operations over structured values, MolADT makes representational assumptions explicit, supports deterministic validation and localized transformations, and provides hooks for symmetry-aware and Bayesian workflows. We provide a reference implementation in Haskell (open-source, archived with DOI) and worked examples demonstrating delocalised/multicentre bonding, validation invariants, reaction extensions, and group actions relevant to geometric learning.
Scientific Contribution: We (1) introduce a representation-level framework that treats molecular representations as well-defined syntactic contracts rather than serializations, (2) formalize a layered typed IR capturing constitution/geometry/annotations, and (3) provide an open reference implementation intended to enable more controlled and interpretable benchmarking of molecular ML pipelines.
@@ 1 Introduction
The choice of molecular representation affects cheminformatics workflows such as molecular property prediction and generative design [1]. Different representations can enable the exploration of different classes of molecules, as well as relationships between a molecule’s chemistry and it’s properties. Different representations can also make it easier to structure the information required by machine learning models. The mathematical primitive used in a representation, and data type (i.e. the “format” of a representaion on a computer) can both constrain and illuminate workflows of a cheminformatician [2].
Because representation choices can change what is easy to learn, what is considered “valid” and what constitutes a meaningful perturbation, model benchmarks are only comparable when the representational substrate is specified and stable. Our goal is not to report new state-of-the-art prediction metrics; rather, we propose a representation designed to make representational assumptions explicit and therefore benchmarkable.
We propose treating molecular representation as a typed intermediate representation: a structured value with explicit invariants and well-scoped edit operations, rather than as a surface syntax (SMILES/SELFIES) whose syntax are enforced only by external parsers/normalisers. Benchmarking ML in chemistry is undermined by a hidden confounder: the molecular representation is often treated as a neutral “input format”, but it actually defines the invariants, and edit operations that models learn over. A typed intermediate representation makes the syntax explicit, so benchmarking can be done on a stable substrate rather than on brittle serializations.
Here we explore representation of molecules via Algebraic Data Types (ADTs): composite data types within a programming language, formed from simpler constructors for atoms, connectivity, stereochemistry, and geometry, together with composible operations on that structure.
After decades of work, string encodings such as SMILES [3] and, SELFIES [4] remain widely used for interchange and as model inputs. They are compact and interoperable, but they primarily serialize a 2D graph-level description and therefore do not directly carry the 3D information (coordinates, conformers) that many tasks ultimately depend on [5, 6]. Coverage of stereochemistry and edge cases varies across toolchains, and complex bonding patterns (general delocalisation, multicentre bonding, many organometallic conventions) are typically handled via additional conventions or specialized tokens rather than a uniform semantic model. For machine learning, a second set of issues matters: token sequences can be syntactically invalid, small token edits can induce large structural changes, and symmetry/invariance structure is not represented as a first-class object. These factors can force models to spend capacity on learning a surface grammar rather than chemistry, complicating controlled benchmarking and interpretation.
Although SMILES provides a convenient, widely adopted linearization of molecular graphs, it is a highly constrained language: only a small subset of character sequences correspond to syntactically valid, parseable SMILES, and an even smaller subset correspond to chemically meaningful molecules. As a result, any approach that “samples SMILES strings” in the unconstrained sense is dominated by invalid outputs. This is not merely an efficiency nuisance; it fundamentally confounds empirical comparisons, because performance becomes driven by how often a method stumbles into valid syntax rather than by its ability to model molecular structure. For this reason, throughout we treat syntactic validity as part of the modeling/inference problem rather than an after-the-fact filter, and we avoid evaluations that implicitly reward or penalize methods based on arbitrary amounts of invalid-string rejection.
Rather than extending a string grammar until it behaves like a language, as proposed in [1], one can treat the molecule itself as a first‑class program object in an ordinary programming language, with explicit types and operations defined on it. In a functional paradigm, ADTs paired with pure functions may provide composability, equational reasoning about transformations on molecules, and immutability, helping to make invariants and exceptions explicit. We propose that if molecules are to be generated, validated, optimised and transformed by programs, representing them directly as typed data – rather than as tokens that must be decoded – offers a more transparent and verifiable route.
In this work, we represent molecules as ADTs so that parsing, validation, and transformation operate on structured molecular objects (atoms, bonds, bonding systems, geometry, and annotations).
@@
A brief note on terminology. We distinguish (i) storage or transmission formats that serialise molecules for storage and exchange (e.g., SMILES, InChI, SDF) from (ii) internal data models (computational data type) used for editing, featurisation, and learning. In practice, these are sometimes discussed under the same umbrella term “representation”. Making the distinction explicit matters for benchmarking: many properties that ML pipelines rely on—validity conditions, canonicalisation conventions, and the definition of “local edits”—are properties of a data model (plus its validators), not of a character string in isolation.
Functional programming has been used in cheminformatics primarily as an implementation choice for existing workflows (e.g., leveraging immutability, strong typing, and safer composition) [9, 10]. Our focus is different: we propose a representation design in which the molecular data model itself is explicitly typed and layered, so that invariants and transformations become part of the representation’s semantics rather than being enforced indirectly by external parsers and toolkit conventions.
@@
Below are some definitions used, defined informally. For formal definitions, see [11, 12].
Definition 1 (Data type)
A data type is a compile-time contract specifying which values may exist and which operations are meaningful on those values.
Definition 2 (Algebraic Data Types (ADTs))
An Algebraic Data Type (ADT) is a data type whose possible values are defined by constructors (products/records and sums/variants, possibly recursive) and which is deconstructed by pattern matching.
1.1 Molecular Graphs, hypergraphs, and multigraphs
@@ 1.1.1 Molecular Graphs
Many molecular representations are use molecular graphs and thus inherit their strengths and limitations for representation[13]. When decoded, the string based representations such as SMILES and SELFIES correspond to molecular graphs, and the atom and bond “blocks” of SDF molfiles represent a molecular graph, stored as a connection table.
Definition 3 (Graph)
A graph $G=(V,E)$ is a pair consisting of a set of vertices $V$ and an edge set $E\subseteq V\times V$.
Definition 4 (Molecular Graph)
A molecular graph is a graph, $G=(V,E)$, where vertices, $V$, represent atoms and edges, $E\subseteq V\times V$, represent bonds. Molecular graphs are connected, undirected, and labeled. Vertex labels may include information such as element or charge, and edge labels may include the bond type as a categorical label, a bond order, or stereochemical annotations [14, 13].
@@ 1.1.3 Multigraphs
A multigraph $G=(V,E)$ has vertex set $V$ and a multiset (bag) $E$ of edges where edges may repeat.
As explained by Dietz[14] Vertices in a molecular multigraph represent atoms, which can be labeled as with molecular graphs, but edges have a different meaning. Each edge is not a bond, but a bonding system. A single edge between two vertices represents a bond between atoms. Where there are multiple edges between pairs of vertices, each edge represents the atom’s share of electrons in bonds in which the atoms participate. For instance, in benzene, each C-C pair has two edges, one representing the 2c-2e bond between adjacent carbon atoms, and another edge representing the delocalised elections contributed to the ring.
By using a multigraph approach, it is possible to label bonds with information about the binary bonding relationships between atoms, as well as describing the delocalised electrons that persist over a subset of atoms [14].
@@ 1.1.4 Dietz representation
Dietz factors structure into constitution (connectivity and electron sharing), configuration (discrete stereochemical arrangements), and conformation (continuous 3D geometry) [14]. In MolADT, we follow Dietz for constitution and supply geometry separately via an explicit coordinate layer; this separation allows a single constitution to be paired with multiple conformers while keeping constitutional invariants stable.
In MolADT, we adopt Dietz’s constitution layer as the semantic core for bonding (including delocalised and multicentre systems), and we can attach 3D coordinates as a separate layer. This separation supports geometry-free symbolic manipulation when coordinates are absent, while still enabling stereochemical and conformational reasoning when coordinates are present.
Definition 5 (Constitution)
@@
Intuition before the examples: a “bonding system” is an electron pool shared across one or more atom–atom edges. When the pool spans a single edge, it behaves like an ordinary localised bond. When the pool spans several edges, it encodes delocalisation while keeping all pairwise adjacencies explicit. The following examples illustrate both cases:
@@
The bonding-system representation treats delocalisation as primitive and allows familiar quantities to be derived when needed (e.g., for back-compatibility with bond-order-based tooling). For $b=(s,F)$ we write $n(b)=s$ and $p(b)=|F|$. Dietz defines the electron count associated with an atom $x$ and derived formal charge/bond-order quantities as functions of the bonding systems incident to x. In MolADT these are treated as derived (computed) properties rather than part of the core representation.
Induced labeled multigraph. For some algorithms it is convenient to work with an induced multigraph view: create an edge $(\{i,j\},b)$ for each bonding system $b=(s,F)$ and each $\{i,j\}\in F$, labeled by (id(b), s). This makes explicit how a constitution $(V,B)$ yields a multigraph plus metadata grouping edges into shared-electron systems, preserving pairwise adjacency while encoding delocalisation via edge groupings.
@@
Figure 1 contrasts a Kekulé depiction with a delocalised depiction of benzene. In Dietz’s formulation, the $\pi$ electrons of the benzene ring are represented as a single bonding system (one electron pool) spanning the six C–C ring edges, separate from the $\sigma$ framework.
@@
1.1.5 Beyond benzene: multi-centre and organometallic bonding
To show that the same constitutional machinery extends beyond typical organic examples, we include an organometallic case. A major advantage of the Dietz representation is its general model for bonding, making it straightforward to describe complex systems such as organometallics (e.g., ferrocene) and electron-deficient bonding (e.g., diborane). It is straightforward to describe complex bonding systems such as organometallics, with ferrocene and diborane already given as motivating examples in the original article. Delocalised bonding is represented explicitly via bonding systems rather than through special atom types or bond labels. This stands in contrast to string-based representations such as SMILES, whose grammar encodes aromaticity through atom typing (e.g. c, n, [nH]) and a collection of special-case rules for different heteroaromatic environments.
One possible Dietz-style constitution for ferrocene can be written schematically as
@@ 1.1.5 Beyond benzene: multi-centre and organometallic bonding
where the final three bonding systems respectively capture the $\pi$-electron delocalisation within each cyclopentadienyl ring and an illustrative Fe–Cp back-donation pool. The precise choice of bonding systems is not unique, but the representation makes explicit which interactions are localised and which are genuinely multi-centre and delocalised.
Figure 2: Three-dimensional view of ferrocene showing the metallocene “sandwich” geometry. Reproduced from [17].
@@
The Dietz representation avoids this false dichotomy by separating pairwise adjacency from electron delocalisation. Explicit two-centre edges (Fe–C, C–C, C–H, etc.) are retained to preserve neighbourhood structure and geometry, while delocalised interactions are encoded via labelled bonding systems (s,F)(s,F)—electron pools of size $s$ spanning a finite set of edges $F$. This permits a faithful representation in which Cp aromaticity, Fe–C coordination, and back-donation coexist without inventing artificial bond types or sacrificing chemical semantics.
Listings A.2 and A.3 provide concrete stress-tests of this approach. Diborane contains two $\mathrm{B\text{–}H\text{–}B}$ bridges that are standardly described as two-electron three-centre (3c–2e) bonds, while ferrocene is characterised by $\eta^{5}$ (haptic) coordination between the iron centre and each Cp ring.[18, 19] In both cases, the localised $\sigma$ framework is expressed as an ordinary adjacency set (localBonds), and each bridge or coordination interaction is represented as an explicit bonding system with a specified electron count distributed over multiple edges. Multi-centre bonding is therefore expressed directly, rather than being squeezed into an ill-fitting 2-centre formalism.
Despite satisfying the three requirements identified by Krenn et al. for representations beyond organic chemistry—namely support for delocalised bonding, explicit terminal hydrogens, and advanced stereochemical phenomena such as cumulenes—Dietz-style representations are described in that work (apparently mischaracterised as hypergraphs) as leading to “complicated nested sets of brackets that may be hard to comprehend.”[1] While the set-theoretic notation may be unfamiliar, the underlying constitution of a molecule admits a compact and unambiguous grammar. This stands in contrast to SMILES and SELFIES strings, whose linear grammars rely on overloaded symbols and traversal-dependent conventions to encode bonding semantics.[1]
These limitations are especially acute for organometallic and multi-centre systems. SMILES provides only pairwise bond primitives with a small fixed vocabulary; as a result, such molecules are typically approximated via disconnected charged fragments or dialect-specific conventions (including “zero-order” or haptic workarounds), and encodings need not round-trip or canonicalise consistently across toolchains.[19, 20] SELFIES inherits this restriction because it offers a robust grammar over the same underlying atom–bond graph—guaranteeing syntactic validity but not introducing a native notion of multi-centre electron pools or hapticity.[4, 21]
@@ 1.2.1 SMILES
SMILES encodes a labelled 2D molecular graph as an ASCII string using atoms, bonds, branches, and ring closures, with optional isomeric annotations for stereochemistry. As an interchange format it is compact and widely interoperable, but it does not natively include 3D coordinates or conformational ensembles, and several behaviors depend on toolchain conventions (e.g., aromaticity perception, kekulisation, and canonicalisation) [13, 15, 1, 22, 23, 24, 25]. These factors matter for benchmarking because models trained or evaluated on SMILES inherit representation-specific notions of validity and locality [26]. In practice, limitations arise around (i) resonance/tautomerism and delocalisation (which require multiple structures or conventions), (ii) the breadth of stereochemical cases that can be expressed at the graph level, and (iii) variation across toolchains (e.g., aromaticity perception, canonicalisation).[15] These are active areas of improvement: OpenSMILES provides a community spec, and newer proposals such as BALSA offer a formally specified subset with explicit syntax to reduce ambiguity across implementations [27].
@@
SMILES representations are non-unique: a single molecule can be expressed by multiple valid strings [3]. As well as complicating database searches [15], machine learning algorithms must discern the equality of different representations. Although some machine learning approaches have attempted to utilize multiple SMILES for the same molecule in their training sets and have claimed improvements to their models as a result [34, 35], these gains may reflect a data-augmentation/regularisation effect: the model is trained to be invariant to multiple serializations of the same underlying graph, but they also underline that the learning objective must spend capacity on the string grammar rather than chemistry. This motivates representations whose locality and invariants are explicit. In other words, it exposes a weakness of SMILES for machine learning, not an advantage.
SMILES does not explicitly represent molecular symmetries (graph automorphisms) as first-class objects. For example, the Cyclopentadienyl anion in Figure  3 can be written as [cH-]1cccc1, but the rotational symmetry evident in Fig. 3 is not encoded in the string. Moreover, because SMILES is non-unique, alternative strings such as 1cccc1[cH-] represent the same structure without any explicit indication of equivalence at the representation level.